# 2DMamba: Efficient State Space Model for Image Representation with Applications on Giga-Pixel Whole Slide Image Classification

### v2

@@ 1 Introduction
Beyond applications in general vision tasks, Mamba finds a great potential in computational pathology, particularly for the classification of Giga-pixel Whole Slide Images (WSIs), known as the gold standard for cancer diagnosis [19, 51, 39, 31, 32, 33]. WSIs are high-resolution images of tissue samples, often reaching up to $000\times 100,$ 000 × 100 , 000 pixels at 40x magnification, making them extremely large and rich in spatial detail. Due to their enormous size, WSIs are typically analyzed in a Multiple Instance Learning (MIL) manner: conventional bag-based MIL methods convert a WSI into a “bag” of instances (patches) that are usually aggregated independently, neglecting the spatial awareness among patches [21, 29, 48]. In contrast, Mamba-based methods treat the WSI as a sequence of patches [13, 46], enabling more effective information aggregation and potentially enhancing diagnostic insights. However, they still flatten 2D images into 1D sequences, and spatial discrepancies persist, as illustrated in Fig. 1. Given that cells interact with each other in a coordinated manner across all directions, scanning in multiple directions [26, 40] does not accurately model the complexity of cell-to-cell interactions.
@@ 2 Related work
Figure 2: Left: The overall architecture of 2DMambaMIL for WSI representation. An input WSI is first tiled into patches and these patches are embedded by a feature extractor into a 2D features map. Non-tissue regions are padded with the learnable token to maintain the 2D spatial relationships. The 2D feature map is then fed to $U$ layers of 2D-Mamba blocks, where the key difference, compared with vanilla Mamba block, is our 2D selective scan module. Right: Our 2D selective scan algorithm. It performs parallel horizontal scan and parallel vertical scan for each state dimension $d$ independently. Parameter $C$ then aggregates $N$ state dimensions into a single dimension output $y$.
@@ 3.2 Architecture of 2DMambaMIL
By doing so, 2DMamba aggregates information without spatial discrepancy. In comparison, the hidden state of the vanilla Mamba on a flattened image is given by $h^{1D}_{i}=\sum_{i^{\prime}\leq i}\bar{A}^{\bm{i-i^{\prime}}}\bar{B}x_{i^{\prime}}$ where $i$ denotes the 1D index. The order $i$-$i^{\prime}$ represents the distance between $i,i^{\prime}$ in a flattened sequence, where a higher order (larger distance) may lead to forgetting [42]. This mathematically encapsulates the concept of “spatial discrepancy”. In contrast, 2DMamba achieves the formulation in Eq.(7), where the order $i$-$i^{\prime}$+$j$-$j^{\prime}$ corresponds to the Manhattan distance between $(i^{\prime},j^{\prime})$ and $(i,j)$, thereby preserving the 2D structure. This distance represents a path from $(i^{\prime},j^{\prime})$ to $(i,j)$ that moves horizontally to the right and then vertically downward. This mathematically encapsulates the concept of “spatial continuity”. For instance, the last hidden state for a 3x3 feature map can be expressed as:
@@
Figure 3: Our hardware-aware 2D selective scan operator with efficient caching mechanism and high parallelism. Orange color represents operations on SRAM and green color represents those on HBM. (a) The 1D Mamba scan operator intakes a flattened sequence on HBM. It tiles the input into sub-sequences. Each sub-sequence is loaded from HBM to SRAM, scanned and reduced across $N$ intermediate dimensions, and then written back to HBM. The total memory access complexity is $\mathcal{O}(L)$. (b) A Naive 2D scan operator tiles the 2D feature map by rows and columns, and performs 1D Mamba scans on each row, column, and on $N$ independent state dimensions. This will explicitly instantiate $N$ intermediate feature maps on HBM, resulting in a memory access complexity of $\mathcal{O}(NL)$. (c) Our 2D scan operator tiles the feature map into 2D grids and scans each grid in 2 directions. Intermediate features are reduced inside each tile, only the aggregated result is stored back to HBM. The memory complexity is $\mathcal{O}(L)$. (d) GPU memory hierarchy: SRAMs are small but fast; HBMs have large capacities but are slow. (e) NVIDIA’s CUB BlockScan only supports 1D sequences, with sizes of multiples of $32$. Scanning a two-row grid requires two sequential kernel launches and padding elements. (f) Our SegmentedBlockScan enables scanning multiple rows and columns in parallel. It reduces the amount of memory transactions and padding data.
@@ 3.4 Hardware-Aware 2D Selective Scan
$Acc$ $F1$ $AUC$ $Acc$ $F1$ $AUC$ $Acc$ $F1$ $AUC$ $Acc$ $F1$ $AUC$ $Acc$ $F1$ $AUC$
@@ 4.1 Dataset
We assess 2DMambaMIL on 5 public pathology classification datasets, TCGA-BRCA [1], BRACS [3], PANDA [4], DHMC [50], TCGA-NSCLC and 5 public survival datasets, TCGA-(KIRC, KIRP, LUAD, STAD, UCEC). These datasets cover a variety of organs including breast, prostate, lung, kidney, stomach, and uterine. The number of slides ranges from $261$ to $10614$ and we use $20$x magnification for all these datasets. Details of the datasets are listed in the Supp. D. Following [26], we evaluate our 2DMamba on two natural image datasets of ImageNet-1K classification and ADE20K semantic segmentation.
@@ 4.2 Results
Method #Param. FLOPs mIoU(SS) mIoU(MS)
@@ Acknowledgement
This work was partially supported by USA NSF grants IIS-2123920 [D.S], IIS-2212046 [D.S], IIS-1715985 [H.Q], IIS-1812606 [H.Q], the Canadian Cancer Society Breakthrough Grant [V.Q.H.T], FRQS-CRS-J1 [V.Q.H.T], NSERC-DG RGPIN-2022-05378 [M.S.H] and Amazon Research Award [M.S.H].
@@ B Details of the hardware-aware 2D selective scan operator
State dimension $N$. Tile size $T$.
Input independent $A:(N)$, $D$ and $bias$;
@@
    # loop for $N$ state dimensions
    for d = 1 to $N$ do
@@ C Mathematical derivations of 2DMamba
[…]
@@
[…]
@@
[…]
@@ G Additional ablation studies
$Acc$ $F1$ $AUC$ $Acc$ $F1$ $AUC$ $Acc$ $F1$ $AUC$ $Acc$ $F1$ $AUC$ $Acc$ $F1$ $AUC$
@@
In 2DMamba, $\bar{A}$ is reused for both horizontal and vertical scan to maintain the same number of parameters as the vanilla Mamba [15]. Although we lack theoretical guarantees of optimality, our ablation studies in Tab. S3 demonstrate that 2DMambaMIL with independent $\bar{A}$ achieves comparable performance (within ±$\pm{1\%}$) to 2DMambaMIL with reused $\bar{A}$ across five datasets.
@@
State dimension $N$ PANDA TCGA-BRCA
$Acc$ $AUC$ $Acc$ $AUC$
@@
Table S9: Ablation study on the state dimension $N$ on the PANDA and TCGA-BRCA dataset.
@@ H Standard derivation
$Acc$ $F1$ $AUC$ $Acc$ $F1$ $AUC$ $Acc$ $F1$ $AUC$ $Acc$ $F1$ $AUC$ $Acc$ $F1$ $AUC$

### v3

@@ 1 Introduction
Beyond applications in general vision tasks, Mamba finds a great potential in computational pathology, particularly for the classification of Giga-pixel Whole Slide Images (WSIs), known as the gold standard for cancer diagnosis [18, 51, 38, 31, 32, 30]. WSIs are high-resolution images of tissue samples, often reaching up to $100,000\times 100,000$ pixels at 40x magnification, making them extremely large and rich in spatial detail. Due to their enormous size, WSIs are typically analyzed in a Multiple Instance Learning (MIL) manner: conventional bag-based MIL methods convert a WSI into a “bag” of instances (patches) that are usually aggregated independently, neglecting the spatial awareness among patches [20, 28, 48]. In contrast, Mamba-based methods treat the WSI as a sequence of patches [12, 46], enabling more effective information aggregation and potentially enhancing diagnostic insights. However, they still flatten 2D images into 1D sequences, and spatial discrepancies persist, as illustrated in Fig. 1. Given that cells interact with each other in a coordinated manner across all directions, scanning in multiple directions [25, 39] does not accurately model the complexity of cell-to-cell interactions.
@@ 2 Related work
Figure 2: Left: The overall architecture of 2DMambaMIL for WSI representation. An input WSI is first tiled into patches and these patches are embedded by a feature extractor into a 2D features map. Non-tissue regions are padded with the learnable token to maintain the 2D spatial relationships. The 2D feature map is then fed to $U$ layers of 2D-Mamba blocks, where the key difference, compared with vanilla Mamba block, is our 2D selective scan module. Right: Our 2D selective scan algorithm. It performs parallel horizontal scan and parallel vertical scan for each state dimension $d$ independently. Parameter $C$ then aggregates NN state dimensions into a single dimension output $y$.
@@ 3.2 Architecture of 2DMambaMIL
By doing so, 2DMamba aggregates information without spatial discrepancy. In comparison, the hidden state of the vanilla Mamba on a flattened image is given by $h^{1D}_{i}=\sum_{i^{\prime}\leq i}\bar{A}^{\bm{i-i^{\prime}}}\bar{B}x_{i^{\prime}}$ where $i$ denotes the 1D index. The order $ii^{\prime}$ represents the distance between $i,i^{\prime}$ in a flattened sequence, where a higher order (larger distance) may lead to forgetting [41]. This mathematically encapsulates the concept of “spatial discrepancy”. In contrast, 2DMamba achieves the formulation in Eq.(7), where the order $ii^{\prime}$$jj^{\prime}$ corresponds to the Manhattan distance between $(i^{\prime},j^{\prime})$ and $(i,j)$, thereby preserving the 2D structure. This distance represents a path from $(i^{\prime},j^{\prime})$ to $(i,j)$ that moves horizontally to the right and then vertically downward. This mathematically encapsulates the concept of “spatial continuity”. For instance, the last hidden state for a 3x3 feature map can be expressed as:
$\begin{aligned} h^{1D}_{3,3}=&\bar{A}^{{\color[rgb]{1,0.49,0}\bm{8}}}\bar{B}x_{1,1}+\bar{A}^{{\color[rgb]{1,0.49,0}\bm{7}}}\bar{B}x_{1,2}+\bar{A}^{{\color[rgb]{1,0.49,0}\bm{6}}}\bar{B}x_{1,3}+&h^{2D}_{3,3}=&\bar{A}^{{\color[rgb]{1,0.49,0}\bm{4}}}\bar{B}x_{1,1}+\bar{A}^{{\color[rgb]{1,0.49,0}\bm{3}}}\bar{B}x_{1,2}+\bar{A}^{{\color[rgb]{1,0.49,0}\bm{2}}}\bar{B}x_{1,3}+\\ &\bar{A}^{{\color[rgb]{1,0.49,0}\bm{5}}}\bar{B}x_{2,1}+\bar{A}^{{\color[rgb]{1,0.49,0}\bm{4}}}\bar{B}x_{2,2}+\bar{A}^{{\color[rgb]{1,0.49,0}\bm{3}}}\bar{B}x_{2,3}+&&\bar{A}^{{\color[rgb]{1,0.49,0}\bm{3}}}\bar{B}x_{2,1}+\bar{A}^{{\color[rgb]{1,0.49,0}\bm{2}}}\bar{B}x_{2,2}+\bar{A}^{{\color[rgb]{1,0.49,0}\bm{1}}}\bar{B}x_{2,3}+\\ &\bar{A}^{{\color[rgb]{1,0.49,0}\bm{2}}}\bar{B}x_{3,1}+\bar{A}^{{\color[rgb]{1,0.49,0}\bm{1}}}\bar{B}x_{3,2}+\bar{A}^{{\color[rgb]{1,0.49,0}\bm{0}}}\bar{B}x_{3,3}&&\bar{A}^{{\color[rgb]{1,0.49,0}\bm{2}}}\bar{B}x_{3,1}+\bar{A}^{{\color[rgb]{1,0.49,0}\bm{1}}}\bar{B}x_{3,2}+\bar{A}^{{\color[rgb]{1,0.49,0}\bm{0}}}\bar{B}x_{3,3}\end{aligned}$
@@
Figure 3: Our hardware-aware 2D selective scan operator with efficient caching mechanism and high parallelism. Orange color represents operations on SRAM and green color represents those on HBM. (a) The 1D Mamba scan operator intakes a flattened sequence on HBM. It tiles the input into sub-sequences. Each sub-sequence is loaded from HBM to SRAM, scanned and reduced across NN intermediate dimensions, and then written back to HBM. The total memory access complexity is $\mathcal{O}(L)$. (b) A Naive 2D scan operator tiles the 2D feature map by rows and columns, and performs 1D Mamba scans on each row, column, and on NN independent state dimensions. This will explicitly instantiate NN intermediate feature maps on HBM, resulting in a memory access complexity of $\mathcal{O}(NL)$. (c) Our 2D scan operator tiles the feature map into 2D grids and scans each grid in 2 directions. Intermediate features are reduced inside each tile, only the aggregated result is stored back to HBM. The memory complexity is $\mathcal{O}(L)$. (d) GPU memory hierarchy: SRAMs are small but fast; HBMs have large capacities but are slow. (e) NVIDIA’s CUB BlockScan only supports 1D sequences, with sizes of multiples of $32$. Scanning a two-row grid requires two sequential kernel launches and padding elements. (f) Our SegmentedBlockScan enables scanning multiple rows and columns in parallel. It reduces the amount of memory transactions and padding data.
@@ 3.4 Hardware-Aware 2D Selective Scan
AccAcc $F1$ AUCAUC AccAcc $F1$ AUCAUC AccAcc $F1$ AUCAUC AccAcc $F1$ AUCAUC AccAcc $F1$ AUCAUC
@@ 4.1 Dataset
We assess 2DMambaMIL on 5 public pathology classification datasets, TCGA-BRCA [42], BRACS [2], PANDA [3], DHMC [50], TCGA-NSCLC and 5 public survival datasets, TCGA-(KIRC, KIRP, LUAD, STAD, UCEC). These datasets cover a variety of organs including breast, prostate, lung, kidney, stomach, and uterine. The number of slides ranges from $261$ to $10614$ and we use 2020x magnification for all these datasets. Details of the datasets are listed in the Supp. D. Following [25], we evaluate our 2DMamba on two natural image datasets of ImageNet-1K classification and ADE20K semantic segmentation.
@@ 4.2 Results
Method #Param. FLOPs mIoU($S$) mIoU(MS)
@@ Acknowledgement
This work was partially supported by USA NSF grants IIS-2123920 [D.S], IIS-2212046 [D.S], IIS-1715985 [H.Q], IIS-1812606 [H.Q], the Canadian Cancer Society Breakthrough Grant [V.Q.H.T], FRQS-CRS-J1 [V.Q.H.T], NSERC-DG RGPIN‐2022‐05378 [M.S.H] and Amazon Research Award [M.S.H]. The computational pathology experiments reported in this paper were executed on Korea University servers administered by Professor Jin Tae Kwak.
@@ B Details of the hardware-aware 2D selective scan operator
State dimension NN. Tile size $T$.
Input independent $A:(N)$, $D$ and biasbias;
@@
  # loop for NN state dimensions
  for d = 1 to NN do
@@ C Mathematical derivations of 2DMamba
$h^{hor}_{i,j}$ $=\sum_{j^{\prime}\leq j}\bar{A}^{(j-j^{\prime})}\bar{B}x_{i,j^{\prime}}\ \text{(The vanilla Mamba)}$ (11)
@@
$=\sum_{i^{\prime}\leq i}\bar{A}^{(i-i^{\prime})}\bar{B^{\prime}}h^{hor}_{i^{\prime},j}\ \text{(The vanilla Mamba)}$ (14)
@@
$=\sum_{i^{\prime}\leq i}\sum_{j^{\prime}\leq j}\bar{A}^{(i-i^{\prime}+j-j^{\prime})}\bar{B}x_{i^{\prime},j^{\prime}}\ \text{(Eq.(7))}$ (16)
@@ G Additional ablation studies
AccAcc $F1$ AUCAUC AccAcc $F1$ AUCAUC AccAcc $F1$ AUCAUC AccAcc $F1$ AUCAUC AccAcc $F1$ AUCAUC
@@
In 2DMamba, $\bar{A}$ is reused for both horizontal and vertical scan to maintain the same number of parameters as the vanilla Mamba [14]. Although we lack theoretical guarantees of optimality, our ablation studies in Tab. S3 demonstrate that 2DMambaMIL with independent $\bar{A}$ achieves comparable performance (within $\pm{1\%}$) to 2DMambaMIL with reused $\bar{A}$ across five datasets.
@@
State dimension NN PANDA TCGA-BRCA
AccAcc AUCAUC AccAcc AUCAUC
@@
Table S9: Ablation study on the state dimension NN on the PANDA and TCGA-BRCA dataset.
@@ H Standard derivation
AccAcc $F1$ AUCAUC AccAcc $F1$ AUCAUC AccAcc $F1$ AUCAUC AccAcc $F1$ AUCAUC AccAcc $F1$ AUCAUC